A point moves so that the ratio of its distances from 3x + 4y + 8 = 0 and 4x + 3y – 6 = 0 is 2. Find the equation of its locus.


1
Expert's answer
2021-11-08T16:09:37-0500

A point moves so that the ratio of its distances from "3x + 4y + 8 = 0" and "4x + 3y \u2013 6 = 0" is "2". Let us find the equation of its locus. Let a point "(x,y)" belongs to the locus. Then

"\frac{|3x + 4y + 8|}{\sqrt{3^2+4^2}}: \frac{|4x + 3y – 6|}{\sqrt{4^2+3^2}}=2."

It follows that

"|3x + 4y + 8|=2|4x + 3y \u2013 6|"

which is equivalent to

"|3x + 4y + 8|^2=4|4x + 3y \u2013 6|^2."

Then

"9x^2+16y^2+64+24xy+48x+64y=4(16x^2+9y^2+36+24xy-48x-36y)."

We conclude that

"55x^2+20y^2+72xy-144x-80y+80=0"

is the equation po its locus.